# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ 4.3.2. Data Extraction and Analysis
We present our results in a structured format to ensure clarity. For each research question, we first summarize the insights from the SLR, followed by key takeaways from the survey study. Next, we analyze broader trust literature to contextualize the trust concepts even further. We then synthesize insights from all three sources to identify gaps, similarities, and key takeaways. Finally, we outline opportunities for future research, highlighting areas that require further exploration.
@@ 5.1. Phase1: SLR Results
In our corpus, only 27.8% ($n=5$ out of 18) of the reviewed papers explicitly defined trust, with all definitions borrowed from established concepts rather than introducing new ones. The two primary definitions of trust utilized in the corpus are:
  • “An attitude that an agent will achieve an individual’s goal in a situation characterized by uncertainty and vulnerability” (Lee and See, 2004a). This definition was cited by (Cheng et al., 2023; Wang et al., 2024c; Widder et al., 2021b; Wang et al., 2024c) (80%, $n=4$ of all the papers that defined trust).
  • ”The extent to which a user is confident in and willing to act based on, the recommendations, actions, and decisions of an artificially intelligent decision aid” (Madsen and Gregor, 2000). This definition appeared in  (Weisz et al., 2021) (20%, $n=1$).
Both definitions originate from social science and have been adapted to the Human-Machine context. These definitions emphasize that trust is rooted in the trustee’s vulnerable position, arising from uncertainty about whether the LLM will meet expectations and deliver desired outcomes. This inherent uncertainty motivates the trustee to place trust in the model (Vereschak et al., 2021). Trust is also characterized by the trustee’s positive expectation of the tool’s ability and intention to achieve its goals. Furthermore, trust is described as an attitude, how a person views and evaluates a particular situation, and is not a direct determinant of behavior. While trust can guide behaviors such as reliance on the LLM, this process is subject to individual, organizational, and cultural contexts (Liao and Sundar, 2022; Wang et al., 2024d).
Trust plays a crucial role in shaping the adoption and continued use of LLMs and software development tools (Brown et al., 2024). However, trust alone is insufficient; fostering ”appropriate trust” is essential to ensure that users rely on the model in ways that align with its actual capabilities (Kästner et al., 2021; Cheng et al., 2023; Wang et al., 2024c). Trust is an attitude stakeholders have towards the model, yet it does not necessarily indicate whether it is reliable. For instance, if a user’s trust level does not match the actual soundness of the system. In that case, they might either accept a flawed response or unjustly dismiss accurate output. Neither scenario is desirable.
One possible mediator between trust and appropriate trust is to design for model trustworthiness (Kästner et al., 2021). In the context of SE, a model is considered trustworthy if and only if it functions correctly and it is justified for stakeholders to trust the system (Kästner et al., 2021). This approach can help bridge the gap between user perception and system reliability, ultimately leading to more effective and responsive use of LLMs in SE practices.
@@ 5.2. Phase2: Survey Study Results
We analyzed 25 responses to open-ended questions on definitions of trust in LLMs for SE. One of the authors iteratively coded the survey responses, following an inductive approach to identify recurring themes. The other authors collaboratively verified and refined the final coding to ensure consistency and reliability. This analysis resulted in the following categories:
@@
  • Functional Correctness and Performance: This was the most frequently mentioned theme across all downstream tasks. The participants emphasized the importance of LLMs generating code that meets the specified requirements, passes the test cases, and improves performance. For example, a participant noted: ”Fundamentally, I think that trust in this sense means that the code, test, or summary produced is accurate. That is if I produce a code snippet, the model hasn’t hallucinated any modules and the snippet will accomplish the task specified in the prompt…”. This theme was particularly significant for test case generation, with 75% of participants referencing it in their responses.
@@
  • Understandability: Understandability emerged as an essential attribute, reflecting participants’ desire for outputs that are clear and easy to comprehend without extensive verification. This was especially emphasized for code generation tasks, where 26.32% of responses mentioned it. One participant remarked ”I would consider trust to mean that I would be comfortable including the LLM-generate code into a code review packet without feeling the need to verify the code in advance.”
@@
  • Security and System Integrity: Participants expressed concern regarding the generated code’s security. They emphasized the critical role of ensuring that LLMs do not compromise the systems they generate code for. They underscored the importance of minimizing bugs and vulnerabilities. For instance, one of the participants defined trust in LLMs for code generation as ”Trust is relying on LLMs do not intentionally try to undermine systems they generate code for.” Interestingly, no participants mentioned security and system integrity as essential factors when defining trust in LLMs for test case generation.
@@
  • Usability and Knowledge Enhancement: Participants valued the use of the generated code directly without careful evaluations and the potential of LLMs to provide additional knowledge. One of them stated, ”Trust in LLMS means I can use the code from LLMS directly in my study or work occasion.” Interestingly, the practitioner was a novice with minimal experience in SE and LLMs, reinforcing the importance of educating practitioners to foster calibrated trust in LLMs for SE applications.
@@
  • Reliability and Consistency: This category reflects the emphasis participants placed on long-term performance and consistent behavior of LLMs in SE tasks. It was the least used theme across definitions. One participant stated, ”Trust in LLMs for Code Generation could be interpreted as the percentage of accurate code generated within an extended time frame of use.”
@@
Our study reveals widespread use of LLMs in SE, with 99.8% of participants using them at least weekly. Despite the widespread use, only a moderate proportion trust the models (52.9% for code generation, 75% for test case generation, 54.5% for program repair). Additionally, while 84% believe that conceptualizing trust in LLM4SE is essential, 60% of participants either struggle to differentiate between trust and trustworthiness or fail to recognize the distinction altogether. This highlights the complexity of these concepts and their varying interpretations in public perception.
@@ 5.3. Complementary Analysis for Definition of Trust
Trust is a widely studied concept across disciplines such as HCI, DL, and automation. In AI research, two dominant perspectives on trust exist: (1) trust as a subjective human attitude and (2) trustworthiness as an intrinsic system property. The first perspective, common in HCI, defines trust as a user’s perception of an AI system’s reliability, shaped by factors such as interpretability, confidence, and perceived competence under conditions of vulnerability  (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022; Huang et al., 2024; Ferrario and Loi, 2022; Perrig et al., 2023). In SE literature, trust is often framed similarly as an expectation about an LLM’s ability to function under uncertainty.
The second perspective, more common in DL and AI governance literature, views trustworthiness as an inherent system property, defined by attributes such as robustness, fairness, and security  (Kaur et al., 2022; Li et al., 2023b; Liu et al., 2021). In ML and AI auditing, trustworthiness is associated with fairness metrics and bias mitigation  (Lassiter and Fleischmann, 2024), while in automation, trust calibration ensures that user reliance aligns with actual system performance  (Lee and See, 2004a). Furthermore, some studies argue that trust is a multifaceted and context-dependent concept, cautioning against overly broad or narrow definitions that may limit its practical applicability  (Butler et al., 2004).
Despite these established perspectives, SE literature rarely distinguishes between trust and trustworthiness, leading to conceptual ambiguity. Our analysis also reveals that distrust is often overlooked, despite being a critical component of trust dynamics.
@@ 5.4. Discussion
Our findings reveal a significant gap between how trust is defined in the literature and how practitioners conceptualize it in real-world SE tasks. The review showed that most papers either lacked explicit definitions of trust or borrowed definitions from broader AI and social science research. These definitions describe trust as an attitude toward uncertainty and risk, focusing on the trustee’s perception rather than the system’s actual capabilities. However, our survey study demonstrates that practitioners define trust based on technical factors such as functional correctness, security, and usability, highlighting a fundamental disconnect between research and practice.
To bridge this gap, we propose the following definitions based on our systematic literature review, survey study, and insights from broader trust research. Table 1 shows overview of our definitions.
@@
    Trust Trust in LLM4SE refers to a practitioner’s willingness to rely on the recommendations or outputs of an LLM system for software engineering tasks, based on the belief that the system is capable, reliable, and aligned with their objectives under conditions of uncertainty. This trust is shaped by the system’s perceived and demonstrated performance, including correctness, security, and usability. This definition is supported by broader trust literature, where trust is conceptualized as a subjective attitude shaped by user perception  (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022), as well as our SLR findings, where trust is often framed as expectation-driven. Additionally, our survey results reveal that practitioners define trust more concretely in terms of functional correctness, interpretability, and usability, emphasizing practical reliability over abstract expectations.
@@
    Distrust Distrust in LLM4SE is not merely the absence of trust but signifies a practitioner’s active expectation of risk, failure, or harm in the system’s outputs for a given SE task. This arises when the model is perceived as unreliable, unpredictable, or misaligned with ethical and functional goals, including security vulnerabilities and biased or unsafe recommendations. HCI research highlights that distrust can be driven by inconsistencies in AI behavior, misalignment with user needs, or previous negative experiences  (Degachi et al., 2024). Similarly, our survey results indicate that practitioners actively distrust LLMs when they exhibit frequent inaccuracies or security vulnerabilities, particularly in safety-critical SE tasks. The distinction between trust and distrust is thus crucial, as distrust involves an explicit expectation of failure rather than mere uncertainty.
@@
    Trustworthiness Trustworthiness in LLM4SE is the intrinsic quality of the LLM system that justifies a user’s trust, based on its ability to consistently perform tasks accurately, transparently, securely, and ethically. Unlike trust, which is user-dependent, trustworthiness is an objective property that can be evaluated through measurable factors such as robustness, fairness, security, and reliability. This aligns with DL research, where trustworthiness is defined through system properties such as fairness, robustness, and security  (Kaur et al., 2022; Li et al., 2023b; Zhang et al., 2024). In contrast to trust, which is a user-driven perception, trustworthiness represents an objective measure of AI system reliability. Our SLR found that trustworthiness in SE is rarely explicitly defined, leading to confusion between perceived and actual system reliability. Survey participants also struggled with this distinction, reinforcing the need to establish trustworthiness as a measurable property rather than an assumed user sentiment.
@@
One key insight from our survey is that trust in LLM4SE is highly task-dependent. While 75% of respondents expressed trust in LLMs for test case generation, trust was lower for code generation (52.9%) and program repair (54.5%). This finding aligns with prior HCI research, which suggests that trust is dynamic and varies based on task complexity and perceived risk  (Jakesch et al., 2019). It also supports observations in DL and automation research, where trust is shaped by domain-specific considerations  (Lee and See, 2004a).
@@
Additionally, our results indicate that practitioners often equate trust with reliability. Many survey participants described trust as something that builds over time through consistent correctness of outputs, rather than as an abstract attitude toward AI systems. This perspective is consistent with findings from automation and robotics, where trust is calibrated through ongoing system performance rather than being static  (Lee and See, 2004a). Similarly, in SE, trust in tools is linked to their reputation, verification, and ability to reduce the need for manual oversight  (Hou and Jansen, 2022; Rutinowski et al., 2024a). Our results suggest that trust in LLM4SE follows a similar pattern, reinforcing the need to incorporate both subjective and system-oriented factors in trust research.
@@
Finally, security considerations emerged as a crucial component of how practitioners define trust, yet none of the reviewed SLR papers explicitly incorporated security into trust definitions. Several survey respondents associated trust with the model’s ability to produce safe, non-malicious code and prevent vulnerabilities. This gap highlights the need to integrate security into trustworthiness discussions, as emphasized in DL research on robust and privacy-preserving AI (Zhang et al., 2024; Li et al., 2023b). Additionally, institutional trust perspectives propose that trust in AI should extend beyond individual system performance to include regulatory oversight and transparency (Knowles and Richards, 2021). These insights could inform how trust in LLM4SE is conceptualized, particularly as AI governance frameworks evolve.
@@ 5.5. Opportunities for Future Work
Our findings suggest several critical avenues for future research. One key area that needs attention is the refinement and validation of trust-related concepts. The definitions that we have proposed are grounded in insights from literature review and practitioner perspectives. However, their applicability in real-world SE tasks requires further empirical assessment. A promising direction for future research could involve conducting longitudinal studies to examine how trust evolves as practitioners interact with LLM for SE tasks. Additionally, task-specific experiments focussing on different SE tasks, such as code generation, test case generation, and program repair, could help clarify whether trust in LLMs is highly task-dependent or whether users develop overarching heuristics that generalize across diverse SE contexts.
@@
Another important area of future research is the study of distrust in the context of LLM4SE. While trust and trustworthiness have been fairly studied, distrust in LLM4SE remains an underexplored phenomenon. Our findings suggest that distrust is not merely the absence of trust but an active expectation of failure, risk, or potential harm. Future research should investigate how distrust develops and whether it leads to outright rejection of LLM-generated outputs or selective verification based on task complexity. Additionally, designing interventions to mitigate distrust—such as enhanced transparency, explicit failure explanations, and confidence estimations—could play a crucial role in improving user confidence in LLMs. Understanding the dynamics between trust, distrust, and corrective user behavior will be essential for designing AI-assisted development environments that align with practitioner expectations.
@@
Lastly, there is a need to further explore the distinction between trust and trustworthiness. Our findings suggest that this distinction is not well understood by practitioners, as 60% of survey respondents struggled to differentiate between the two concepts. By exploring the interactions between trust, trustworthiness, and related constructs such as confidence, belief, and reliability, researchers could contribute to more precise conceptual frameworks for trust in AI. This would help bridge the gap between theoretical understandings of trust and the practical considerations that shape how users interact with LLMs in SE contexts.
@@
Summary of $RQ_{1}$: Definition of Trust Concepts in LLM4SE Our study found that only 27.8% ($n=5$ out of 18) of reviewed papers explicitly defined trust, borrowing established definitions from social science and AI research rather than introducing new ones. Survey results from 25 participants revealed that practitioners conceptualize trust in LLMs based on functional correctness, security, understandability, and usability, whereas literature primarily frames it as an attitude toward uncertainty and risk. This discrepancy highlights a significant gap between theoretical definitions and real-world expectations. Furthermore, 60% of participants struggled to distinguish between trust and trustworthiness, indicating conceptual confusion. Trust in LLMs was found to be task-dependent, with higher trust levels for test case generation (75%) than for code generation (52.9%) and program repair (54.5%). We propose distinct definitions: trust as a practitioner’s willingness to rely on LLMs based on perceived reliability, distrust as an expectation of failure, and trustworthiness as an intrinsic system property ensuring reliability, security, and ethical alignment. Addressing these differences is crucial for fostering appropriate trust in LLMs for software engineering.

### v2

@@ 4.3.2. Data Extraction and Analysis
We present our results in a structured format to ensure clarity. For each research question, we first summarize the insights from the SLR, followed by key takeaways from the survey study. Next, we analyze the broader trust literature to contextualize trust concepts even further. We then synthesize insights from all three sources to identify gaps, similarities, and key takeaways. Finally, we outline opportunities for future research, highlighting areas that require further exploration.
@@ 5.1. Phase1: SLR Results
In our corpus, only 27.8% ($n=5$ out of 18) of the reviewed articles explicitly defined trust, with all definitions borrowed from established concepts rather than introducing new ones (Cheng et al., 2023; Wang et al., 2024c; Widder et al., 2021; Borg, 2024; Weisz et al., 2021). The two primary definitions of trust utilized in the corpus are the following:
  • ”An attitude that an agent will achieve an individual’s goal in a situation characterized by uncertainty and vulnerability” (Lee and See, 2004). This definition was cited by (Cheng et al., 2023; Wang et al., 2024c; Widder et al., 2021; Borg, 2024) ($n=4$ of all articles that defined trust).
  • ”The extent to which a user is confident and willing to act based on the recommendations, actions, and decisions of an artificially intelligent decision aid” (Madsen and Gregor, 2000). This definition appeared in (Weisz et al., 2021) ($n=1$).
Both definitions originate from social science and have been adapted to the Human-Machine context. These definitions emphasize that trust is rooted in the trustee’s vulnerable position, arising from uncertainty about whether the LLM will meet expectations and deliver desired outcomes. This inherent uncertainty motivates the trustee to place trust in the model (Vereschak et al., 2021). Trust is also characterized by the trustee’s positive expectation of the tool’s ability and intention to achieve its goals. Furthermore, trust is described as an attitude, as how a person views and evaluates a particular situation, and is not a direct determinant of behavior. Although trust can guide behaviors such as LLM trust, this process is subject to individual, organizational, and cultural contexts (Liao and Sundar, 2022; Wang et al., 2024c).
Trust plays a crucial role in shaping the adoption and continued use of LLMs and software development tools (Brown et al., 2024). However, trust alone is insufficient; fostering ”appropriate trust” is essential to ensure that users rely on the model in ways that align with its actual capabilities (Kästner et al., 2021; Cheng et al., 2023; Wang et al., 2024c). Trust is an attitude that stakeholders have towards the model, yet it does not necessarily indicate whether it is reliable. For instance, if a user’s trust level does not match the actual soundness of the system. In that case, they might either accept a flawed response or unjustly dismiss accurate output. Neither scenario is desirable.
One possible mediator between trust and appropriate trust is the design of trustworthiness models (Kästner et al., 2021). In the context of SE, a model is considered trustworthy if and only if it functions correctly and it is justified for stakeholders to trust the system (Kästner et al., 2021). This approach can help bridge the gap between user perception and system reliability, ultimately leading to more effective and responsive use of LLMs in SE practices.
@@ 5.2. Phase2: Survey Study Results
We analyzed 25 responses to open-ended questions on the definitions of trust in LLMs for SE. One of the authors iteratively coded the survey responses, following an inductive approach to identify recurring themes. The other authors collaboratively verified and refined the final coding to ensure consistency and reliability. This analysis resulted in the following categories:
@@
  • Functional Correctness and Performance: This was the most frequently mentioned theme across all downstream tasks. The participants emphasized the importance of LLMs generating code that meets the specified requirements, passes the test cases, and improves performance. For example, a participant noted: ”Fundamentally, I think trust in this sense means that the code, test, or summary produced is accurate. That is if I produce a code snippet, the model hasn’t hallucinated any modules and the snippet will accomplish the task specified in the prompt…”. This theme was particularly significant for the generation of test cases, and 75% of the participants referred to it in their responses.
@@
  • Understandability: Understandability emerged as an essential attribute, reflecting the desire of the participants for a result that is clear and easy to understand without extensive verification. This was especially emphasized for code generation tasks, where 26.32% of the responses mentioned it. One participant remarked ”I would consider trust to mean that I would be comfortable including the LLM-generate code into a code review packet without feeling the need to verify the code in advance.”
@@
  • Security and System Integrity: Participants expressed concern regarding the generated code’s security. They emphasized the critical role of ensuring that LLMs do not compromise the systems they generate code for. They underscored the importance of minimizing bugs and vulnerabilities. For example, one of the participants defined trust in LLMs for code generation as ”Trust is relying on LLMs do not intentionally try to undermine systems they generate code for.” Interestingly, no participants mentioned security and system integrity as essential factors when defining trust in LLMs for test case generation.
@@
  • Usability and Knowledge Enhancement: Participants valued the use of the generated code directly without careful evaluations and the potential of LLMs to provide additional knowledge. One of them stated, ”Trust in LLMS means that I can use the LLMS code directly in my study or work occasion.” Interestingly, the practitioner was a novice with minimal experience in SE and LLMs, reinforcing the importance of educating practitioners to foster calibrated trust in LLMs for SE applications.
@@
  • Reliability and Consistency: This category reflects the emphasis the participants placed on long-term performance and consistent behavior of LLMs in SE tasks. It was the least used theme in all the definitions. One participant stated, ”Trust in LLMs for Code Generation could be interpreted as the percentage of accurate code generated within an extended time frame of use.”
@@
Our study reveals widespread use of LLMs in SE, with 99.8% of participants using them at least weekly. Despite widespread use, only a moderate proportion trust the models (52.9% for code generation, 75% for test case generation, 54.5% for program repair). Furthermore, while 84% believe that conceptualizing trust in LLM4SE is essential, 60% of the participants struggle to differentiate between trust and trustworthiness or fail to recognize the distinction. This highlights the complexity of these concepts and their varying interpretations in public perception.
@@ 5.3. Complementary Analysis for Definition of Trust
Trust is a widely studied concept across disciplines such as HCI, DL, and automation. In AI research, two dominant perspectives on trust exist: (1) trust as a subjective human attitude and (2) trustworthiness as an intrinsic system property. The first perspective, common in HCI, defines trust as a user’s perception of the reliability of an AI system, shaped by factors such as interpretability, confidence, and perceived competence under conditions of vulnerability (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022; Huang et al., 2024; Ferrario and Loi, 2022; Perrig et al., 2023). In the SE literature, trust is often framed in a similar way to an expectation about the ability of an LLM to function under uncertainty (Ghofrani et al., 2022; Paulus et al., 2013).
The second perspective, more common in the DL, views trustworthiness as an inherent property of the system, defined by attributes such as robustness, fairness, and security (Kaur et al., 2022; Li et al., 2023b; Liu et al., 2021). In AI audits, trustworthiness is associated with fairness metrics and bias mitigation (Lassiter and Fleischmann, 2024), while in automation, trust calibration ensures that user reliance aligns with actual system performance (Lee and See, 2004). Furthermore, some studies argue that trust is a multifaceted and context-dependent concept, cautioning against overly broad or narrow definitions that may limit its practical applicability (Butler et al., 2004).
Despite these established perspectives, the SE literature rarely distinguishes between trust and trustworthiness, leading to conceptual ambiguity. Our analysis also reveals that distrust is often overlooked, despite being a critical component of trust dynamics.
@@ 5.4. Discussion
Our findings reveal a disconnect between research and practice: the literature defines trust abstractly as a perception of risk , while practitioners define it using concrete technical factors like functional correctness, security, and usability.
To bridge this gap, we propose the following definitions based on our systematic review of the literature, survey study, and insights from broader trust research. Table 2 shows an overview of our definitions.
@@
    Trust Trust in LLM4SE refers to the willingness of a practitioner to rely on the recommendations or outputs of an LLM system for software engineering tasks, based on the belief that the system is capable, reliable, and aligned with their objectives under conditions of uncertainty. This trust is shaped by the perceived and demonstrated performance of the system, including correctness, security, and usability. This definition is supported by broader trust literature, where trust is conceptualized as a subjective attitude shaped by user perception (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022), as well as our SLR findings, where trust is often framed as expectation driven. Furthermore, the results of our survey reveal that practitioners define trust more concretely in terms of functional correctness, interpretability, and usability, focusing on practical reliability over abstract expectations.
@@
    Distrust Distrust in LLM4SE is not merely the absence of trust, but signifies practitioners’ active expectation of risk, failure, or harm in the system output for a given SE task. This occurs when the model is perceived to be unreliable, unpredictable, or misaligned with ethical and functional goals, including security vulnerabilities and biased or unsafe recommendations. HCI research highlights that distrust can be driven by inconsistencies in AI behavior, misalignment with user needs, or previous negative experiences (Degachi et al., 2024). Similarly, our survey results indicate that professionals actively distrust LLMs when they exhibit frequent inaccuracies or security vulnerabilities, particularly in safety-critical SE tasks. The distinction between trust and distrust is thus crucial, as distrust involves an explicit expectation of failure rather than mere uncertainty.
@@
    Trustworthiness The trustworthiness in LLM4SE is the intrinsic quality of the LLM system that justifies a user’s trust based on its ability to consistently perform tasks accurately, transparently, securely, and ethically. Unlike trust, which is user-dependent, trustworthiness is an objective property that can be evaluated through measurable factors such as robustness, fairness, security, and reliability. This aligns with the DL research, where trustworthiness is defined through system properties such as fairness, robustness, and security (Kaur et al., 2022; Li et al., 2023b; Zhang et al., 2024). In contrast to trust, which is a user-driven perception, trustworthiness represents an objective measure of the reliability of the AI system. Our SLR found that trustworthiness in SE is rarely explicitly defined, leading to confusion between perceived and actual system reliability. Survey participants also struggled with this distinction, reinforcing the need to establish trustworthiness as a measurable property rather than an assumed user sentiment.
@@
One key insight from our survey is that trust in LLM4SE is highly task-dependent. While 75% of respondents expressed trust in LLMs for test case generation, trust was lower for code generation (52.9%) and program repair (54.5%). This finding is consistent with previous HCI research, which suggests that trust is dynamic and varies based on task complexity and perceived risk (Jakesch et al., 2019). It also supports observations in DL and automation research, where trust is shaped by domain-specific considerations (Lee and See, 2004).
@@
Furthermore, our results indicate that practitioners often equate trust with reliability. Many survey participants described trust as something that builds over time through consistent correctness of outputs, rather than as an abstract attitude toward AI systems. This perspective is consistent with findings from Automation and Robotics, where trust is calibrated through ongoing system performance rather than static (Lee and See, 2004). Similarly, in SE, the trust in the tools is related to their reputation, verification, and the ability to reduce the need for manual supervision (Hou and Jansen, 2022; Rutinowski et al., 2024). Our results suggest that trust in LLM4SE follows a similar pattern, reinforcing the need to incorporate subjective and system-oriented factors in trust research.
@@
Finally, security considerations emerged as a crucial component of how practitioners define trust, yet none of the reviewed SLR papers explicitly incorporated security into trust definitions. Several survey respondents associated trust with the model’s ability to produce safe and non-malicious code and prevent vulnerabilities. This gap highlights the need to integrate security into trustworthiness discussions, as emphasized in DL research on robust and privacy-preserving AI (Zhang et al., 2024; Li et al., 2023b). Furthermore, institutional trust perspectives propose that trust in AI should extend beyond individual system performance to include regulatory oversight and transparency (Knowles and Richards, 2021). These insights could inform how trust in LLM4SE is conceptualized, particularly as AI governance frameworks evolve.
@@ 5.5. Opportunities for Future Work
Our findings suggest several critical avenues for future research. One key area that needs attention is the refinement and validation of concepts related to trust. The definitions that we have proposed are based on insights from the literature review and the perspective of practitioners. However, their applicability in real-world SE tasks requires further empirical assessment. A promising direction for future research could involve conducting longitudinal studies to examine how trust evolves as practitioners interact with LLM for SE tasks. Additionally, task-specific experiments focused on different SE tasks, such as code generation, test case generation, and program repair, could help clarify whether trust in LLMs is highly task dependent or whether users develop overarching heuristics that generalize across diverse SE contexts.
While our corpus has derived a definition of trust from the social sciences, trust is a topic well-studied across a variety of fields. Psychological safety, for example, is similar to trust in our context, as it refers to an individual’s perception that they can take interpersonal risks without fear of negative consequences (Edmondson, 1999). Although this concept aligns closely with our definition of trust, previous papers in LLM4SE have largely overlooked it. Future studies should broaden the exploration of trust by drawing inspiration from other domains.
@@
Another important area of future research is the study of distrust in the context of LLM4SE. Although trust and trustworthiness have been fairly studied, distrust in LLM4SE remains an underexplored phenomenon. Our findings suggest that distrust is not simply the absence of trust, but an active expectation of failure, risk, or potential harm. Future research should investigate how distrust develops and whether it leads to outright rejection of LLM-generated outputs or selective verification based on task complexity. In addition, designing interventions to mitigate distrust, such as improved transparency, explicit explanations of failures, and confidence estimates, could play a crucial role in improving user confidence in LLM. Understanding the dynamics between trust, distrust, and corrective user behavior will be essential for designing AI-assisted development environments that align with practitioner expectations.
@@
Lastly, there is a need to explore the distinction between trust and trustworthiness further. Our findings suggest that this distinction is not well understood by practitioners, as 60% of the survey respondents struggled to differentiate between the two concepts. By exploring the interactions between trust, trustworthiness, and related constructs such as confidence, belief, and reliability, researchers could contribute to more precise conceptual frameworks for trust in AI. This would help bridge the gap between theoretical understandings of trust and the practical considerations that shape how users interact with LLMs in SE contexts.
@@
Summary of $RQ_{1}$: Definition of Trust Concepts in LLM4SE Our study found that only 27.8% ($n=5$ out of 18) of reviewed papers explicitly defined trust, borrowing established definitions from social science and AI research rather than introducing new ones. The survey revealed that practitioners conceptualize trust in LLMs based on functional correctness, security, understandability, and usability, while the literature primarily frames it as an attitude toward uncertainty and risk. This discrepancy highlights a significant gap between theoretical definitions and real-world expectations. Furthermore, 60% of the participants struggled to distinguish between trust and trustworthiness, indicating conceptual confusion. The trust in LLMs was found to depend on tasks, with higher trust levels for test case generation (75%) than for code generation (52.9%) and program repair (54.5%). We propose distinct definitions: trust as the willingness of a practitioner to rely on LLMs based on perceived reliability, distrust as an expectation of failure, and trustworthiness as an intrinsic property of the system that ensures reliability, security, and ethical alignment. Addressing these differences is crucial for fostering the appropriate trust in LLMs for software engineering.